# An Online Learning Approach to Prompt-based Selection of Generative Models

### v2

@@ Abstract
Selecting a sample generation scheme from multiple text-based generative models is typically addressed by choosing the model that maximizes an averaged evaluation score. However, this score-based selection overlooks the possibility that different models achieve the best generation performance for different types of text prompts. An online identification of the best generation model for various input prompts can reduce the costs associated with querying sub-optimal models. In this work, we explore the possibility of varying rankings of text-based generative models for different text prompts and propose an online learning framework to predict the best data generation model for a given input prompt. The proposed PAK-UCB algorithm addresses a contextual bandit (CB) setting with shared context variables across the arms, utilizing the generated data to update kernel-based functions that predict the score of each model available for unseen text prompts. Additionally, we leverage random Fourier features (RFF) to accelerate the online learning process of PAK-UCB and establish a $\widetilde{\mathcal{O}}(\sqrt{T})$ regret bound for the proposed RFF-based CB algorithm over $T$ iterations. Our numerical experiments on real and simulated text-to-image and image-to-text generative models show that RFF-UCB performs successfully in identifying the best generation model across different sample types.
@@ 1 Introduction
Text-based generative artificial intelligence (AI) has found numerous applications in various engineering tasks. A prompt-based generative AI represents a conditional generative model that produces samples given an input text prompt. Over the past few years, several frameworks using diffusion models and generative adversarial networks have been proposed to perform text-guided sample generation tasks for various data domains including image, audio, and video (Reed et al.,, 2016; Pan et al.,, 2018; Xu et al.,, 2018; Ding et al.,, 2021; Singer et al.,, 2022; Huang et al.,, 2023; Podell et al.,, 2024). The multiplicity of developed prompt-based models has led to significant interest in developing evaluation mechanisms to rank the existing models and find the best generation scheme. To address this task, several evaluation metrics have been proposed to quantify the fidelity and relevance of samples created by prompt-based generative models, such as CLIPScore (Hessel et al.,, 2021) and PickScore (Kirstain et al.,, 2023).
Example 1 Stable Diffusion v1.5 PixArt-$\bm{\alpha}$
@@
subfigureExample 1: Stable Diffusion v1.5 attains a higher CLIPScore in generating prompts with term “car” (36.10 versus 35.68) while underperforms for prompts with term “dog” (36.37 versus 37.24).
Example 2 UniDiffuser Stable Diffusion v1.5
Prompts of Type “elephant”
Refer to caption Refer to caption
avg. CLIPScore $\bm{36.67}\ (\pm 0.05)$ $35.08\ (\pm 0.06)$
Prompts of Type “fire hydrant”
Refer to caption Refer to caption
avg. CLIPScore $35.11\ (\pm 0.14)$ $\bm{37.23}\ (\pm 0.05)$
subfigureExample 2: UniDiffuser attains a higher CLIPScore in generating prompts with the term “elephant” (36.67 versus 35.08) while underperforms for prompts with the term “fire hydrant” (35.11 versus 37.23).
@@
Figure 1: Standard text-to-image (T2I) models exhibit different rankings based on CLIPScore (Hessel et al.,, 2021) across the text prompts.
@@
The existing model selection methodologies commonly aim to identify the generative model with the highest relevance score, producing samples that correlate the most with input text prompts. A well-known example is the CLIPScore for image generation models, measuring the expected alignment between the input text and output image of the model using the CLIP embedding (Radford et al., 2021a, ). While the best-model identification strategy has been frequently utilized in generative AI applications, this approach does not consider the possibility that the involved models can perform differently across text prompts. However, it is possible that one model outperforms another model in responding to text prompts from certain categories, while that model performs worse in generating samples for other text categories. Figure 1 shows such examples where two standard text-to-image models exhibit different rankings across the text prompts with terms “dog”/“car” and “elephant”/“fire hydrant”. In general, the different training sets and model architectures utilized to train text-based models can result in the models’ varying performance in response to different text prompts, which is an important consideration in the selection of text-based generative models for various text prompts.
@@
In this work, we aim to develop a learning algorithm to identify the best generative model for a given input prompt, using observed prompt/generated samples collected from the models in the previous sample generation iterations. Since the goal of such text-based model selection is to minimize the data queries from suboptimal generative models for an input text prompt, we view the model selection task as an online learning problem, where after each data generation the learner updates the prediction on which generative model performs the best in response to different text prompts. Here, the goal of the online learner is to utilize the previously generated samples to accurately guess the generation model with the best performance for the incoming text prompt. An optimal online model selection method will result in a bounded regret value, measured in comparison to the sample generation from the groundtruth-best model for the text prompts.
@@
The described online learning task can be viewed as a contextual bandit (CB) problem studied in the bandit literature (Langford and Zhang,, 2007; Li et al.,, 2010). In the CB task, the online learner observes a single context variable (the text prompt in our setting) and predicts the best arm for the current input context. Specifically, we focus on the kernel-based method to predict the score of each available model for an incoming prompt. As the text prompt is shared across the models, the CB task may be suboptimally addressed by the standard linear CB (Chu et al.,, 2011) and kernelized CB (Valko et al.,, 2013) algorithms in the literature, as these algorithms train a single model to predict the reward functions for all the arms. To relax this constraint, we propose Per-Arm Kernelized UCB (PAK-UCB) to address the online prompt-based selection of generative models. According to the PAK-UCB approach, the learner utilizes UCB-scores from arm-specific kernel-based prediction functions to choose the generative model for the incoming text prompt and subsequently update the kernel-based prediction rule based on the generated data for the upcoming iterations. We prove that the proposed PAK-UCB achieves an $\widetilde{O}(\sqrt{T})$ regret bound over a horizon of $T$ iterations.
@@
Since the user applying the CB-based model selection approach may have limited compute power and not be able to afford growing computational costs in the online learning process, we propose to utilize the random Fourier features (RFF) framework (Rahimi and Recht, 2007a, ) to balance the computational load between the iterations of PAK-UCB. We discuss that in the proposed PAK-UCB, the computational cost per iteration will grow cubically as $O(t^{3})$ with iteration $t$. To address the growing cost per iteration of kernel-UCB, we leverage the RFF approach and develop the proxy RFF-UCB algorithm which approximates the solution to PAK-UCB, while the computational costs grow only linearly $O(t)$. We show that the regret bound for PAK-UCB will approximately hold for RFF-UCB, and therefore, RFF-UCB provides an efficient proxy to the PAK-UCB algorithm, which can be run in devices with lower computation budget.
@@
Finally, we present the results of several numerical experiments to show the efficacy of our proposed PAK-UCB and RFF-UCB in the online selection of conditional generative models. In our experiments, we test several pre-trained and simulated text-to-image and image-captioning (image-to-text) models, where different models lead to different rankings of CLIPScore values across sample types. Our numerical results suggest a fast convergence of the proposed online learning algorithms to the best model available for different prompt types. Moreover, we apply the RFF-UCB method and show how the algorithm can infer the model with higher CLIPScore with a growing accuracy as the iteration grows. In our experiments, the proposed PAK-UCB and RFF-UCB outperform various baseline methods including the standard LinUCB (Chu et al.,, 2011) and KernelUCB (Valko et al.,, 2013) methods. The main contributions of our work can be summarized as:
@@
  • Studying the prompt-based selection of conditional generative models to improve the performance scores over every individual model
@@
  • Developing the contextual bandit-based PAK-UCB and RFF-UCB algorithms for the online selection of prompt-based generative models
@@
  • Providing the theoretical analysis of the regret and computational costs of PAK-UCB and RFF-UCB online learning methods
@@
  • Presenting numerical results on the online selection of generative models based on the incoming prompt using PAK-UCB and RFF-UCB
@@ 2 Related Work
(Automatic) Evaluation of conditional generative models. Evaluating the conditional generative models has been studied extensively in the literature. For text-to-image (T2I) generation, earlier methods primarily rely on the Inception score (Salimans et al.,, 2016) and Fréchet inception distance (Heusel et al.,, 2017). More recent works propose reference-free metrics for robust automatic evaluation of T2I and image captioning, with notable examples being CLIPScore (Hessel et al.,, 2021) and PickScore (Kirstain et al.,, 2023). Kim et al., (2022) propose a mutual-information-based metric, which attains consistency across benchmarks, sample parsimony, and robustness. To provide a holistic evaluation of T2I models, several works focus on multi-objective evaluation. Astolfi et al., (2024) propose to evaluate conditional image generation in terms of prompt-sample consistency, sample diversity, and fidelity. Kannen et al., (2024) introduce a framework to evaluate T2I models regarding cultural awareness and cultural diversity. Masrourisaadat et al., (2024) examine the performance of several T2I models in generating images such as human faces and groups and present a social bias analysis. Another line of study explores evaluation approaches using large language models (LLMs). Tan et al., (2024) develop LLM-based evaluation protocols that focus on the faithfulness and text-image alignment. Peng et al., (2024) introduce a GPT-based benchmark for evaluating personalized image generation. For evaluation of text-to-video (T2V) generation, Huang et al., (2024) introduce VBench as a comprehensive evaluation of T2V models in terms of quality and consistencey.
@@
(Kernelized) Contextual bandits.  The contextual bandits (CB) is an efficient framework for online decision-making with side information (Langford and Zhang,, 2007; Foster et al.,, 2018), which is widely adopted in domains such as recommendation system and online advertisement (Li et al.,, 2010). A key to its formulation is the relationship between the context (vector) and the expected reward. In linear CB, the reward is assumed to be linear to the context vector (Li et al.,, 2010; Chu et al.,, 2011). To incorporate non-linearity, Valko et al., (2013) propose kernelized CB, which assumes the rewards are linear-realizable in a reproducing kernel Hilbert space (RKHS). To address the growing computation, recent work leverages the assumption that the kernel matrix is often approximately low-rank and uses Nyström approximations (Calandriello et al.,, 2019, 2020; Zenati et al.,, 2022).
@@ 3 Preliminaries
3.1 CLIPScore
CLIPScore (Hessel et al.,, 2021) is a widely used automatic metric to evaluate the alignment of text-to-image/video (T2I/V) and image captioning models. Let $(y,x)\in{\mathcal{Y}}\times{\mathcal{X}}$ be any text-image pair. We denote by $\bm{c}_{y}\in{\mathbb{S}}^{d-1}:=\{z\in{\mathbb{R}}^{d}:\|z\|_{2}=1\}$ and $\bm{v}_{x}\in{\mathbb{S}}^{d-1}$ the (normalized) embeddings of text $y\in{\mathcal{Y}}$ and image $x\in{\mathcal{X}}$, respectively, both extracted by CLIP (Radford et al., 2021b, ). The CLIPScore (Hessel et al.,, 2021) is given by

### v3

@@ Abstract
Selecting a sample generation scheme from multiple prompt-based generative models, including large language models (LLMs) and prompt-guided image and video generation models, is typically addressed by choosing the model that maximizes an averaged evaluation score. However, this score-based selection overlooks the possibility that different models achieve the best generation performance for different types of text prompts. An online identification of the best generation model for various input prompts can reduce the costs associated with querying sub-optimal models. In this work, we explore the possibility of varying rankings of text-based generative models for different text prompts and propose an online learning framework to predict the best data generation model for a given input prompt. The proposed PAK-UCB algorithm addresses a contextual bandit (CB) setting with shared context variables across the arms, utilizing the generated data to update kernel-based functions that predict the score of each model available for unseen text prompts. Additionally, we leverage random Fourier features (RFF) to accelerate the online learning process of PAK-UCB. Our numerical experiments on real and simulated text-to-image and image-to-text generative models show that RFF-UCB performs successfully in identifying the best generation model across different sample types. The code is available at: github.com/yannxiaoyanhu/dgm-online-select.
@@ 1 Introduction
Large language models (LLMs) and text-guided generative AI models have found numerous applications in various engineering tasks. A prompt-guided generative AI represents a conditional generative model that produces output samples given an input text prompt. Over the past few years, several frameworks have been developed to train generative models and perform text-guided sample generation in various domains, including text, image, and video (Touvron et al.,, 2023; Gemini-Team et al.,, 2025; OpenAI et al.,, 2024; DeepSeek-AI et al.,, 2025; Rombach et al.,, 2022; Bao et al., 2023b, ; Podell et al.,, 2024; Singer et al.,, 2022; Bai et al.,, 2023; Liu et al.,, 2024). The great number of available LLMs and prompt-guided generative models has led to significant interest in developing evaluation mechanisms to rank the existing models and find the best model from a group of available ones. To address this task, several evaluation metrics have been proposed to quantify the fidelity of samples generated by prompt-based generative models, such as CLIPScore (Hessel et al.,, 2021) and PickScore (Kirstain et al.,, 2023) for prompt-guided image and video generation, and BLEU score (Papineni et al.,, 2002) and BERTscore (Zhang et al.,, 2020) for LLMs.
Example on LLMs Gemini-2.5-Flash Qwen-Plus
Code Completion
Refer to caption Refer to caption
Pass rate (%) $\bm{92.32}\ (\pm 0.43)$ $86.46\ (\pm 0.71)$
Code Translation
Refer to caption Refer to caption
Pass rate (%) $48.78\ (\pm 0.52)$ $\bm{82.26}\ (\pm 1.95)$
subfigureExample 1: Gemini-2.5-Flash attains higher pass rate on (Python) code completion on the HumanEval benchmark ($92.3$% versus $86.5$%) while underperforms for Java-to-C++ translation on the HumanEval-X benchmark ($48.8$% versus $82.3$%).
Example on T2I models Stable Diffusion v1.5 PixArt-$\bm{\alpha}$
@@
subfigureExample 2: Stable Diffusion v1.5 attains a higher CLIPScore in generating MS-COCO prompts with term “car” (36.10 versus 35.68) while underperforms for MS-COCO prompts with term “dog” (36.37 versus 37.24).
@@
Figure 1: Widely-used LLMs and text-to-image models could exhibit different rankings across the text prompts with different categories.
@@
The existing model selection methodologies commonly aim to identify the generative model with the highest averaged fidelity score over the distribution of text prompts, producing samples that, on average, align the most with input text prompts. A well-known example is the averaged CLIPScore for text-to-image models, measuring the expected alignment between the input text and output image of the model using the CLIP embedding (Radford et al., 2021a, ). While the best-averaged score model selection strategy has been frequently utilized in generative AI applications, this approach does not consider the possibility that the involved models can perform differently across text prompts.
However, we highlight the possibility that one model outperforms another model in responding to text prompts from certain categories, while that model performs suboptimally in responding to prompts from other categories. Figure 1 shows examples of both LLMs and text-to-image models, where two widely-used models exhibit different rankings across the text prompts from different categories (code generation and translation tasks for LLMs, and MS-COCO prompts with terms “dog”/“car”). In general, the different training sets and model architectures of LLMs and prompt-guided models can result in the models’ varying performance in response to different text prompts, which is an important consideration in assigning prompts of various types to a group of prompt-input generative models. [^1] The full models’ response codes in the figure’s experiment are provided at: https://github.com/yannxiaoyanhu/dgm-online-select.
Figure 2: Selection of text-to-image (T2I) models based on the CLIPScore: The conventional best-score model selection assigns all prompts to the single model with the best averaged CLIPScore. In contrast, our proposed online approach performs a prompt-level model selection by leveraging the UCB contextual bandit algorithm to assign the model to the prompt, learning from previous data generations.
@@
In this work, we aim to develop a learning algorithm to identify the best generative model for a given input prompt, using observed prompt/generated samples collected from the models in the previous sample generation rounds (Figure 2). Since the goal of prompt-based model selection is to avoid sample generation queries from suboptimal generative models for a given text prompt, we view the model selection task as an online learning problem, where after each data generation, the learner updates the prediction on which generative model performs the best in response to input text prompts. Here, the objective of the online learner is to utilize the previously generated samples to accurately guess the generation model with the best performance for the incoming text prompt. An optimal online model selection method will result in a bounded regret value, measured in comparison to the sample generation from the groundtruth-best model for the text prompts.
@@
We highlight that the described online learning task can be viewed as a contextual bandit (CB) problem widely studied in the bandit literature (Langford and Zhang,, 2007; Li et al.,, 2010; Chu et al.,, 2011; Agrawal and Goyal,, 2013). In the CB task, the online learner observes a single context variable (the text prompt in our setting) and predicts the best arm for the current input context. Specifically, we focus on the kernel-based method to predict the score of each available model for an incoming prompt. As the text prompt is shared across the models, the CB task may be suboptimally addressed by the common-weight formulation of linear CB (Chu et al.,, 2011) and kernelized CB (Valko et al.,, 2013) algorithms, as these implementations of the algorithms learn a single shared weight vector to predict the reward functions for all the arms (i.e., generative models in our setting). To relax this constraint, we propose Per-Arm Kernelized UCB (PAK-UCB) to address the online prompt-based selection of generative models. According to the PAK-UCB approach, the learner utilizes the computed UCB-scores of arm-specific kernel-based prediction functions to choose the generative model for the incoming text prompt and subsequently update the kernel-based prediction rule based on the generated data for the future rounds. We attempt to theoretically analyze the proposed PAK-UCB and show that its variant could achieve a regret bound of $\widetilde{O}(\sqrt{T})$ over $T$ rounds.
@@
Since the user applying the CB-based model selection approach may have limited compute power and not be able to afford growing computational costs in the online learning process, we propose applying the random Fourier features (RFF) framework (Rahimi and Recht, 2007a, ) to reduce the computational load of PAK-UCB. We discuss that in the proposed PAK-UCB, the computational cost per iteration will grow cubically as $O(t^{3})$ with iteration $t$. We demonstrate that after leveraging the RFF method, our developed proxy RFF-UCB algorithm can effectively approximate the solution to PAK-UCB with the computational costs increasing only linearly $O(t)$.
@@
Finally, we present the results of several numerical experiments to show the efficacy of our proposed PAK-UCB and RFF-UCB in the online selection of prompt-guided generative models and LLMs. In our experiments, we test several pre-trained and simulated text-to-image, image-captioning (image-to-text), and language models, where different models could lead to different rankings across various sample types. Our numerical results suggest a fast convergence of the proposed online learning algorithms to the best prompt-based selection of the available models in response to different prompt types. In our experiments, the proposed PAK-UCB and RFF-UCB could perform better than several baseline methods, including the common-weight implementation of LinUCB (Chu et al.,, 2011) and KernelUCB (Valko et al.,, 2013) methods. The following is a summary of this work’s contributions:
@@
  • Studying the prompt-based selection of prompt-guided generative models to improve the prompt-level performance scores,
@@
  • Developing the contextual bandit algorithms of PAK-UCB and RFF-UCB algorithms for the online selection of prompt-based generative models,
@@
  • Theoretically analyzing the regret and computational costs of PAK-UCB and RFF-UCB online learning methods
@@
  • Presenting numerical results on the prompt-based selection of generative models using PAK-UCB and RFF-UCB.
@@ 2 Related Work
(Automatic) Evaluation of conditional generative models. Evaluating the conditional generative models has been studied extensively in the literature. For text-to-image (T2I) generation, earlier methods primarily rely on the Inception score (Salimans et al.,, 2016) and Fréchet inception distance (Heusel et al.,, 2017). More recent works propose reference-free metrics for robust automatic evaluation of T2I and image captioning, with notable examples being CLIPScore (Hessel et al.,, 2021) and PickScore (Kirstain et al.,, 2023). Kim et al., (2022) propose a mutual-information-based metric, which attains consistency across benchmarks, sample parsimony, and robustness. To provide a holistic evaluation of T2I models, several works focus on multi-objective evaluation. Astolfi et al., (2024) propose to evaluate conditional image generation in terms of prompt-sample consistency, sample diversity, and fidelity. Kannen et al., (2024) introduce a framework to evaluate T2I models regarding cultural awareness and cultural diversity. Masrourisaadat et al., (2024) examine the performance of several T2I models in generating images such as human faces and groups and present a social bias analysis. Another line of study explores evaluation approaches using large language models (LLMs). Tan et al., (2024) develop LLM-based evaluation protocols that focus on the faithfulness and text-image alignment. Peng et al., (2024) introduce a GPT-based benchmark for evaluating personalized image generation. In addition, a line of study leverages human feedback for scoring/ranking generated images and improving T2I models (Xu et al.,, 2023). For evaluation of text-to-video (T2V) generation, Huang et al., (2024) introduce VBench as a comprehensive evaluation of T2V models in terms of quality and consistency. Finally, we note the prompt-aware diversity scores, Conditional-Vendi (Jalali et al.,, 2023) and Schur-Complement-Entropy (Ospanov et al., 2024a, ), developed for prompt-guided generative models.
@@
(Kernelized) Contextual bandits.  The contextual bandits (CB) is an efficient framework for online decision-making with side information (Langford and Zhang,, 2007; Foster et al.,, 2018), which is widely adopted in domains such as recommendation system and online advertisement (Li et al.,, 2010) and resource allocations (Lim et al.,, 2024). A key to its formulation is the relationship between the context (vector) and the expected reward. In linear CB, the reward is assumed to be linear to the context vector (Li et al.,, 2010; Chu et al.,, 2011). To incorporate non-linearity, Valko et al., (2013) propose kernelized CB, which assumes the rewards are linear-realizable in a reproducing kernel Hilbert space (RKHS). To address the growing computation, recent work leverages the assumption that the kernel matrix is often approximately low-rank and uses Nyström approximations (Calandriello et al.,, 2019, 2020; Zenati et al.,, 2022).
Prompt-based Selection of LLMs and Generative Models. We note that a line of study proposes offline methods for learning a universal prompt-to-model assignment rule by training a neural network model over a dataset of ranked responses of the models to a large set of prompts (Luo et al.,, 2024; Qin et al.,, 2024; Frick et al.,, 2025). Unlike these methods, our proposed approach follows an online learning strategy which can significantly lower the computational and statistical costs of finding a satisfactory assignment rule. In addition, as we later discuss in the numerical results, the online nature of our proposed algorithm can improve the adaptability of the model to the changes in the user’s prompt distribution and models’ behavior over the iterations of evaluating the models in real time.
Multi-armed bandit selection of generative models. We note that the references (Hu et al., 2025a, ; Rezaei et al.,, 2025) propose multi-armed bandit (MAB) approaches to the selection of unconditional generative models without an input text prompt. Specifically, Hu et al., 2025a propose an online learning framework for selecting generative models and develop the FID-UCB algorithm for the online selection of image-output generative models based on the FID score (Heusel et al.,, 2017). Rezaei et al., (2025) show that a mixture of several unconditional generative models can attain higher diversity scores than each individual model, and propose the Mixture-UCB MAB framework to find the best mixture of the generative models in terms of the kernel distance Bińkowski et al., (2018) and Renyi kernel entropy Jalali et al., (2023); Ospanov et al., 2024b . We note that the mentioned papers study the selection of unconditional generative models, different from our problem setting of selecting the conditional prompt-guided generative models.
@@ 3 Preliminaries
3.1 Kernel Methods and Random Fourier Features
[…]
Every kernel function $k$ will satisfy the positive-definiteness condition, that means $\sum_{i=1}^{n}\sum_{j=1}^{n}c_{i}c_{j}k(y_{i},y_{j})\geq 0$ holds for every integer $n\in{\mathbb{N}}_{+}$, $y_{1},\cdots,y_{n}\in{\mathbb{R}}^{d}$, and $c_{1},\cdots,c_{n}\in{\mathbb{R}}$. Furthermore, we call a kernel function $k$ shift invariant if $k(y,y^{\prime}):=k(y-y^{\prime})$ for every $y,y^{\prime}\in{\mathbb{R}}^{d}$. A well-known example is the radial basis function (RBF) kernel with bandwidth parameter $\sigma>0$:
$k_{\textup{RBF}}(y,y^{\prime})=\exp\Bigl{(}\frac{-\|y-y^{\prime}\|_{2}^{2}}{2\sigma^{2}}\Bigr{)}.$
Kernel ridge regression (KRR). Given empirical labeled samples $(y_{1},s_{1}),\cdots,(y_{n},s_{n})$, where $\{y_{i}\in{\mathbb{R}}^{d}\}_{i=1}^{n}$ are feature variables and $\{s_{i}\in{\mathbb{R}}\}_{i=1}^{n}$ are target variables, respectively, the kernel ridge regression assumes that for a $w^{\star}\in{\mathcal{H}}$ we have $\mathbb{E}[s_{i}|y_{i}]=\phi(y)^{\top}w^{\star}$ for any $i=1,\cdots,n$. For regularization parameter $\alpha\geq 0$, KRR estimator is
$\widehat{s}_{\textup{KRR}}(y):=k_{y}^{\top}(K+\alpha I_{n})^{-1}v$
for every $y\in{\mathbb{R}}^{d}$, where $K=[k(y_{i},y_{j})]_{i,j=1}^{n}\in{\mathbb{R}}^{n\times n}$ denotes the kernel matrix, $v:=[s_{1},\cdots,s_{n}]^{\top}\in{\mathbb{R}}^{n}$, and $k_{y}=[k(y_{1},y),\cdots,k(y_{n},y)]^{\top}\in{\mathbb{R}}^{n}$. The KRR estimator can be interpreted as the solution $w^{\star}$ to the ridge regression:
$\widehat{w}:=\operatorname*{arg\,min}_{w\in{\mathcal{H}}}\sum_{i=1}^{n}\bigl{(}\phi(y_{i})^{\top}w-s_{i}\bigr{)}^{2}+\alpha\|w\|^{2},$
Random Fourier features (RFF). To reduce the computational costs of kernel methods for shift-invariant kernels, Rahimi and Recht, 2007b proposes the random Fourier features framework. Specifically, Bochner’s Theorem (Rudin,, 2017) shows that for every shift-invariant kernel $k(y,y^{\prime})=\kappa(y-y^{\prime})$ satisfying $\kappa(0)=1$, the Fourier transform of $\kappa$ provides a valid probability density function, which we denote by $p\in\Delta({\mathbb{R}}^{d})$, such that $k(y,y^{\prime})=\mathbb{E}_{w\sim p}[e^{iw^{\top}(y-y^{\prime})}],$ with $i$ being the imaginary unit. Following this property, the RFF approach independently samples $w_{1},\cdots,w_{D}\sim p$ and proposes the empirical mean proxy for $k(y,y^{\prime})$ as $\frac{1}{D}\sum_{j=1}^{D}e^{iw_{j}^{\top}(y-y^{\prime})}$. Due to the real Fourier transform of the even function $\kappa$, we can simplify the equation as
$\varphi(y)=\frac{1}{\sqrt{D}}\bigl{[}\cos(w_{1}^{\top}y),\sin(w_{1}^{\top}y),.,\cos(w_{D}^{\top}y),\sin(w_{D}^{\top}y)\bigr{]}$ (1)
[…]
$\widetilde{s}_{\textup{KRR}}(y):=(\widetilde{\Phi}^{*}\widetilde{\Phi}+\alpha I_{2D})^{-1}\widetilde{\Phi}^{*}v,$
[…]
3.2 CLIPScore for Evaluating Text-to-Image Models
CLIPScore (Hessel et al.,, 2021) has been widely-used to evaluate the alignment of samples generated by text-to-image/video (T2I/V) and image captioning models. Let $(y,x)\in{\mathcal{Y}}\times{\mathcal{X}}$ be a text-image pair. We denote by $\bm{c}_{y}\in{\mathbb{S}}^{d-1}:=\{z\in{\mathbb{R}}^{d}:\|z\|_{2}=1\}$ and $\bm{v}_{x}\in{\mathbb{S}}^{d-1}$ the (normalized) embeddings of text $y\in{\mathcal{Y}}$ and image $x\in{\mathcal{X}}$, respectively, both extracted by CLIP (Radford et al., 2021b, ). The CLIPScore (Hessel et al.,, 2021) is given by